Moly is 3 years younger than Havier. In 3 year, havier will be as twice as molly age.
Answer by checkley75(3666)

M=H-3
H+3=2M SUBSTITUTING (H-3)FOR M IN THE SECOND EQUATION WE GET:
H+3=2(H-3)
H+3=2H-6
H-2H-6-3
-H=-9
H=9 HAVIER IS NOW 9 YEARS OLD & MOLLY IS 9-3=6 YEARS OLD.
SO IN THREE YEARS HAVIER WILL BE 9+3=12 OR HAVIER WILL BE TWICE AS OLD AS MOLLY IS.
6*2=12
12=12